In a company an average of 3 out of every 5 staffs take MC on daily basis. A random sample of 10 staffs is selected. Find the probability that Exactly 6 take MC in a particular week.


1
Expert's answer
2021-06-18T04:47:42-0400

Let "X=" the number of staffs taking MC: "X\sim Bin(n, p)."

Given "n=10, p=\dfrac{3}{5}=0.6"


"P(X=6)=\dbinom{10}{6}(0.6)^6(1-0.6)^{10-6}"

"=210(0.6)^6(0.4)^{4}=0.250822656"

The probability that Exactly 6 take MC in a particular week is 0.2508.